# Simul-LLM: A Framework for Exploring High-Quality Simultaneous Translation with Large Language Models

### v1

@@ 1 Introduction
Modern large language models (LLMs) contain at least several billion and up to trillions of parameters and are remarkably capable across a wide range of tasks. Pretrained on humongous amounts of unlabeled data, they have demonstrated incredible emergent capabilities. With minor prompt adjustments, such as including instructions and examples, LLMs are often capable of near state-of-the-art performance set by highly customized solutions. The performance of these models is further enhanced when fine-tuned for specialized downstream tasks, often exceeding the performance of previously cutting-edge solutions. Given their rapidly evolving capabilities, LLMs and their application have become a focused topic of research within NLP academia.
@@
2 Background
While a range of work is relevant to this paper, we will only provide a focused and high level review of large language models, LLMs applied towards machine translation, and simultaneous translation as an application space. Readers interested in additional details should engage further with cited works in these areas.
2.1 Large Language Models
Language modeling via transformers, autoregressive (Irie et al., 2019) or bi-directional (Devlin et al., 2019), has been popular nearly since the architecture’s inception, with bi-directional language modeling via BERT demonstrating particularly potent results for its time. In recent years, however, research on other approaches to this task has largely slowed down in favor of rapidly scaling model parameters and employing massive amounts of high-quality data to train on (Kaplan et al., 2020), resulting in the advent of purely generative Large Language Models (LLMs) like GPT-3 (Brown et al., 2020), LLaMa (Touvron et al., 2023), and others.
At their core, LLMs generate the most likely subsequent token (usually at the sub-word level) given all previous tokens in a sequence. Regardless of this simple, core concept, their size and complexity allows for potent pattern recognition and demonstrates numerous emergent capabilities (e.g. simple math, summarization). Additionally, when fine-tuned for specific tasks, this class of models is capable of near-equivalent performance to specialized solutions while, in many cases, maintaining their generalizability to other tasks. For example, this can allow a single end-to-end model to serve as both a high-quality multilingual translation agent and a chat assistant, opening up interesting, practical deployment opportunities.
@@ 2.2 Large Language Models for Neural Machine Translation
While LLMs are capable of effectively zero-shot sentence-to-sentence neural machine translation (NMT) (Vilar et al., 2023), their performance can still be improved via simple techniques. Prompt construction has been demonstrated to be critical to LLM performance, both before and after fine-tuning Zhang et al. (2023). One-shot or few-shot performance via In-Context Learning (ICL) can produce near competitive results with fine-tuned LLMs for translation and can even be employed to enhance fine-tuned model performance (Vilar et al., 2023; Xu et al., 2023).
One particularly interesting area of study related to LLMs applied towards NMT remains whether or not to fully fine-tune a given model or engage in Parameter-Efficient Fine-Tuning (PEFT) (Mangrulkar et al., 2022). Early work in this area demonstrated some potential for smaller models (Üstün and Cooper Stickland, 2022), and the accessibility that PEFT provides designers in terms of fine-tuning on low-to-mid performance hardware setups renders it desirable. One of the most popular forms of PEFT freezes a LLM’s weights and adds Low-Rank Adaptation (LoRA) (Hu et al., 2022) adapters between layers (other forms of PEFT exist, but adapter-based PEFT is extremely common so we refer to adapter-based PEFT simply as PEFT hereafter). While fully fine-tuned NMT LLMs tend to suffer from some level of catastrophic-forgetting (Kirkpatrick et al., 2017), intuitively, PEFT-based NMT LLMs should not suffer from any loss of off-task performance, as adapters can be loaded or detached depending on whether or not a given user is prompting for a translation. Given all of these factors, PEFT is an attractive option for NMT LLMs.
@@ 2.3 Simultaneous Translation
As a subset of typical NMT, simultaneous translation (SimulMT) focuses on engaging in translation (write decisions) while balancing the amount of available source context (read decisions) to reduce translation latency. This necessarily increases the difficulty of translating a sequence in one language to a sequence in another language, especially when structural and/or syntactical differences exist in the language pair. As a brief example, we can consider translating from a subject-verb-object (SVO) language, such as English, to a subject-object-verb (SOV) language like German. Based on the available source context in English, we may have to guess at the translation in German without access to the necessary context to effectively make that prediction. This is exemplified in Figure 1, where the verb "repaired" must be predicted multiple time-steps before the necessary context is available in the incremental source.
@@
Chunk-wise variations are also possible, where for multiple consecutive translation steps, or write decisions, words or subwords from the last determined most-likely beam are employed to cut down on latency.
@@ 3 Simul-LLM: an Open-Source SimulMT LLM Fine-tuning Framework
SimulMT is an underexplored application space for LLMs, at the moment, and there is plenty of room to improve performance. To facilitate the rapid development of solutions for SimulMT LLMs (fine-tuned for SimulMT) or NMT LLMs (fine-tuned for NMT) adapted for SimulMT, we choose to develop and provide an open-source framework written in PyTorch for researchers to actively employ for future experiments. We call this framework Simul-LLM, and it bridges the gap between the development of fine-tuning agents via popular libraries and proper SimulMT evaluation. Simul-LLM is poised to support both selective classical SimulMT systems as well as fine-tuning and evaluation for a variety of LLM systems such as Falcon (Almazrouei et al., 2023), LLaMa (Touvron et al., 2023) and Mistral-based (Jiang et al., 2023) models. The high-level components of Simul-LLM include a fine-tuning wrapper and a SimulMT evaluation agent for every supported LLM; in case of classical models, only the evaluation agent is supported in-framework.
BitsAndBytes|.3.1 Fine-tuning Wrapper and Features
@@
LLM Support and Extensibility: The proposed Simul-LLM currently supports Falcon and will imminently support Mistral and LLaMa for SimulMT fine-tuning. The fine-tuning wrapper is constructed with a focus on extensibility, allowing for rapid expansion to other LLMs assuming users hold a minimal level of LLM-specific knowledge.
Multiple Prompt Structures: Translation quality, as demonstrated by numerous prior works, varies significantly with prompt structure. As such, the fine-tuning wrapper of Simul-LLM supports multiple kinds of prompt structures, all of which we will validate later in this paper. In the interest of supporting adapting NMT LLMs to SimulMT, Simul-LLM supports NMT fine-tuning in addition to supporting prompt structures that allow for strict wait-k fine-tuning structures.
PEFT and Full Model Fine-tuning: While PEFT (Mangrulkar et al., 2022) has recently become popular, full model fine-tuning is still a valuable option. Simul-LLM supports both PEFT and full model fine-tuning, although PEFT is recommended for most low-to-mid resource hardware setups. We provide recommended PEFT configurations (as seen in Figure 2) for all supported models, and integrate interface for new PEFT configurations when extending to new LLMs.
Flexible Quantization:
@@
library (seen in the Quant Config in Figure 2), which enables flexible fixed-point quantization. For most low-to-mid hardware setups, we recommend quantizing in 4-bit floating point via 4-bit NormalFloat (nf4).
@@
object and a specified response template.
Supervised Fine-tuning Agent:
While custom fine-tuning and optimization loops can be useful, existing and popular options are quite capable. Given that, Simul-LLM’s fine-tuning wrapper surrounds the transformers
supervised fine-tuning trainer and interfaces with it directly. Hyperparameters are fed directly to it as well as PEFT configs, quantization configs, and a specified prompt structure. The trainer itself handles saving model (or LoRA adapter) checkpoints, engaging in validation runs, and optimizing as defined by the provided hyperparameters.
@@ 3.2 Evaluation Agent and Features
Support for Multiple Decoding Strategies: Variable latency constraints for possible inference environments demand flexibility in decoding strategies. As such, evaluation agents for Simul-LLM support several decoding strategies, including greedy and naive decoding, subword-based beam search for single-word decoding, and variations on Speculative Beam Search (SBS) (Zheng et al., 2019), including single word-based SBS and chunk-wise SBS.
Efficient Inference via Custom Generation Stopping Criteria: To avoid excessive generation, Simul-LLM evaluation agents employ custom generation stopping criteria where applicable. This is most relevant in the case of greedy decoding based on singular words, where generation can halt on the first white space delimiter being detected, but constructing more complex criteria is simple given provided examples.
Scoring and Latency via SimulEval: SimulEval (Ma et al., 2020) is the premier SimulMT evaluation framework. Simul-LLM evaluation agents interface seamlessly with SimulEval, which handles the incremental source context and manages translation scoring and latency tracking regarding the translation hypothesis.
@@ 4 Adapting NMT LLMs to SimulMT
To intuit why this is an issue, suppose that an NMT LLM is accustomed to receiving the entire source sequence x before outputting a word or token . If engaging in a low-lagging wait-k where , then the output can now only be based on source context up to . Assuming the NMT LLM would typically rely on some source context where , then it is missing what should be critical information in its translation decision. We have conducted preliminary quantitative studies on this front by employing the proposed Simul-LLM framework, and these results are presented in Section 7.
@@ 5.1 Split Source-Target Prompt Structure
This lack of loss filtering can be especially problematic near the end of a given sequence’s translation. As a simple example, suppose that a given source sequence is of length and . In the first prompt structure where up to source words have been supplied and where , the LLM is effectively being fine-tuned for varying wait-k values ranging from wait-i to wait-k in a single example (as is predicted from to which is wait-k, is predicted from to which is wait-, and so so). If to the point where it is close to normal NMT levels of context, where , then the LLM is no longer being effectively fine-tuned with an appropriate amount of source context for a given translation (write) decision schedule.
@@ 5.2 Single Output Word Prompt Structure
The above problem can be entirely side-stepped by our proposed second approach, single output token prompt structure, that embeds only the current target translation hypothesis within the model output.
As illustrated in the second prompt structure in Figure 4, instead of allowing target translation hypotheses from previous time-steps to be incorporated into the loss, the proposed prompt structure shifts those previous translation hypotheses into the prompt. Combined with the expanded examples that form a rigorous wait-k curriculum in terms of the fine-tuning dataset (rigorous meaning a complete curriculum as opposed to a random subset), inference behavior can be copied exactly for every fine-tuning example, thus completely closing the context mismatch between fine-tuning and inference.
@@ 6 Evaluation Methodology
To validate our proposed solutions to the aforementioned challenges and to test the capabilities of Simul-LLM as the first SimulMT LLM open-source framework, we engaged in several experiments allowing for comparisons among classical non-LLM-based NMT (Vaswani et al., 2017) / SimulMT architectures (Ma et al., 2019), NMT LLMs adapted for SimulMT, and SimulMT LLMs. All mentioned LLMs are fine-tuned Falcon-7B models, but Simul-LLM features an easy to extend framework and we intend to add results for other popular LLMs such as LLaMa-7B and Mistral-7B soon.
@@
SimulMT LLMs with Proposed Prompt | Wait-3 Fine-tuning LLM | 19.99 (3.41) | 23.68 (3.64) |
Wait-7 Fine-tuning LLM | 20.82 (3.44) | 25.18 (3.61) | |
Wait-7 Fine-tuning LLM (k=7) | 23.09 (6.71) | 28.92 (6.87) |
@@ 6.1 Dataset Selection and Preprocessing
No standardized dataset exists for SimulMT with LLMs. Due to its popularity in the speech-to-text simultaneous translation (SimulST) track, we employ MuST-C for our experiments222This allows for future work that explores multi-modal simultaneous LLMs, engaging in SimulST via a cascaded model structure with a transcription model for the source speech or a joint speech/text-to-text framework.. For the purposes of adapting MuST-C for text-to-text usage, we preprocess the dataset and filter out certain acoustic indicators (e.g., floating "-" characters representing pauses). In some cases, this resulted in significant changes to some samples of the test set, such as the removal of (Laughter) and (Applause) acoustic indicators that may take up a good portion of the samples.
@@
6.2 Training, Fine-Tuning, and Evaluation Hyperparameters and Hardware Details
All classical models were trained on two NVIDIA 32GB V100s and validated on a single V100. All LLMs were fine-tuned via PEFT Mangrulkar et al. (2022) on a single NVIDIA 40GB A40 in bfloat16 and evaluated on a single V100 in float32. Simul-LLM seamlessly integrates with SimulEval (Ma et al., 2020) for the purpose of these evaluations. Classical transformer baselines were trained via Fairseq (Ott et al., 2019), an easily extensible sequence to sequence toolkit.
Classical models were trained with typical hyperparameters provided in Fairseq examples. All LLMs were fine-tuned with identical hyperparameters, employing a constant learning rate of 3e-4 and were optimized via Adam with around 4K warmup updates and batch sizes of 40 samples. LoRA adapter parameters were an of 16 and a value of 64, resulting in a total of around 40M added parameters during fine-tuning, with a dropout value of 0.1. For fair comparison, classical models were of a similar, although not quite identical, size in terms of parameter count.
Additionally, all LLMs were fine-tuned while quantized with NormalFloat4 (nf4) quantization. A small performance boost was observed when removing this quantization during inference, so all models did not engage with nf4 quantization during inference. NMT LLMs were fine-tuned for one epoch, as overfitting was observed beyond that point, which we intuit to be possibly due to the well-documented ability of LLMs to quickly memorize training sets (Biderman et al., 2023). In contrast, SimulMT LLMs were fine-tuned for 2M random examples out of 5M examples on the expanded dataset due to computational constraints.
@@ 7.1 Exploration of Adapting NMT LLMs to SimulMT
In Table 1, we provide a breakdown of the performance of several different models, decoding strategies, and wait-k schedules. Regarding our exploration related to adapting NMT LLMs to SimulMT, we also include results related to our implementation of Speculative Beam Search (SBS) (Zheng et al., 2019). As demonstrated by these results, compared with classical models, LLMs fine-tuned for NMT are very capable of SimulMT upon being adapted during inference (even exceeding the score of the classical NMT transformer on en-de that performs non-simultaneous translation). It is worth reiterating that classical architectures typically engage in subword-based wait-k whereas we employ word-based wait-k for LLMs, but the comparisons still serve as a useful reference.

### v4

@@ 1 Introduction
Modern large language models (LLMs) contain at least several billion and up to trillions of parameters and are remarkably capable across a wide range of tasks. Pretrained on humongous amounts of unlabeled data, they have demonstrated incredible emergent capabilities. With minor prompt adjustments, such as including instructions and examples, LLMs are often capable of near state-of-the-art performance set by highly customized solutions. The performance of these models is further enhanced when fine-tuned for specialized downstream tasks, sometimes exceeding the performance of previously cutting-edge solutions. Given their rapidly evolving capabilities, LLMs and their application have become a focused topic of research within NLP academia (Zhao et al., 2023).
@@
2 Background and Motivation
While a range of work is relevant to this paper, we will only provide a focused and high level review of large language models applied towards machine translation and simultaneous translation as an application space. Readers interested in additional details should engage further with cited works in these areas.
@@ 2.1 Large Language Models for Neural Machine Translation
LLMs are capable of effectively zero-shot sentence-to-sentence neural machine translation (NMT) (Vilar et al., 2023), but their performance can still be improved via simple techniques. Prompt construction has been demonstrated to be critical to LLM performance, both before and after fine-tuning Zhang et al. (2023). One-shot or few-shot performance via In-Context Learning (ICL) can produce near competitive results with fine-tuned LLMs for translation and can even be employed to enhance fine-tuned model performance (Vilar et al., 2023; Xu et al., 2023).
One particularly interesting area of study related to LLMs applied towards NMT remains whether or not to fully fine-tune a given model or engage in Parameter-Efficient Fine-Tuning (PEFT) (Mangrulkar et al., 2022). Early work in this area demonstrated some potential for smaller models (Üstün and Cooper Stickland, 2022), and the accessibility that PEFT provides designers in terms of fine-tuning on low-to-mid performance hardware setups renders it desirable. One of the most popular forms of PEFT freezes a LLM’s weights and adds Low-Rank Adaptation (LoRA) (Hu et al., 2022) adapters between layers222Other forms of PEFT exist, but adapter-based PEFT is extremely common so we refer to adapter-based PEFT simply as PEFT hereafter.. While fully fine-tuned NMT LLMs tend to suffer from some level of catastrophic-forgetting (Kirkpatrick et al., 2017), intuitively, PEFT-based NMT LLMs should not suffer from any loss of off-task performance, as adapters can be loaded or detached depending on whether or not a given user is prompting for a translation. Given these factors, PEFT is an attractive option for NMT LLMs.
@@ 2.2 Simultaneous Translation
As a subset of typical NMT, simultaneous translation (SimulMT) focuses on engaging in translation (write decisions) while balancing the amount of available source context (read decisions) to reduce translation lagging behavior/latency. This necessarily increases the difficulty of translating a sequence in one language to a sequence in another language, especially when structural and/or syntactical differences exist in the language pair. As a brief example, we can consider translating from a subject-verb-object (SVO) language, such as English, to a subject-object-verb (SOV) language like German. Based on the available source context in English, we may have to guess at the translation in German without access to the necessary context to effectively make that prediction because of the aforementioned syntactical differences in the language pair.
@@
Chunk-wise variations are also possible, where for multiple consecutive translation steps, or write decisions, words or subwords from the last determined most-likely beam are employed to cut down on computational latency.
2.3 Motivation for Applying LLMs to SimulMT
LLMs are clearly capable of state-of-the-art NMT via a variety of techniques. Extending them towards SimulMT is a natural step, mirroring the progression of classical SimulMT research (i.e. seeking SimulMT improvements from NMT novelties). In this case, smaller language models (LM) have been employed to augment SimulMT efforts before, although usually this augmentation is limited to rendering the translation scheduler more discerning (Indurthi et al., 2022). Instead of just augmenting translation schedulers with smaller, bi-directional LMs, we explore using generative LLMs as direct SimulMT models in this work.
Intuitively, this results in the LLM providing a deeper understanding of source and target languages (via their pretraining on massive, somewhat multilingual corpora) than is typical for most SimulMT translation models. This can matter when encountering context alignment obstacles, where necessary source context is missing for a given translation time-step. In such situations, SimulMT models must infer based on what context does exist and past experiences to determine what the most likely next translated word will be, a task LLMs are naturally suited for.
@@ 3 Simul-LLM: an Open-Source SimulMT LLM Fine-tuning Framework
SimulMT is an underexplored application space for LLMs, at the moment, and there is plenty of room to improve performance. To facilitate the rapid development of solutions for SimulMT LLMs (fine-tuned for SimulMT) or NMT LLMs (fine-tuned for NMT) adapted for SimulMT, we choose to develop and provide an open-source framework written in PyTorch for researchers to actively employ for future experiments. We call this framework Simul-LLM, and it bridges the gap between the development of fine-tuning agents via popular libraries and proper SimulMT evaluation. Simul-LLM is poised to support both selective classical SimulMT systems as well as fine-tuning and evaluation for a variety of LLM systems such as Falcon (Almazrouei et al., 2023), LLaMa (Touvron et al., 2023) and Mistral-based (Jiang et al., 2023) models. The high-level components of Simul-LLM include a fine-tuning wrapper and a SimulMT evaluation agent for every supported LLM; in the case of classical models, only the evaluation agent is supported in-framework.
3.1 Fine-tuning Wrapper and Features
@@
LLM Support and Extensibility: The proposed Simul-LLM currently supports Falcon, Mistral, and Llama-2 fine-tuning (some Llama-2 validation still remains). The fine-tuning wrapper is constructed with a focus on extensibility, allowing for rapid expansion to other LLMs assuming users hold a minimal level of LLM-specific knowledge.
Multiple Prompt Structures: Translation quality, as demonstrated by numerous prior works, varies significantly with prompt structure. As such, the fine-tuning wrapper of Simul-LLM supports multiple kinds of prompt structures, all of which we will validate later in this paper. In the interest of supporting adapting NMT LLMs to SimulMT, Simul-LLM supports NMT fine-tuning in addition to supporting prompts that allow for strict wait-k fine-tuning structures.
BitsAndBytes|.
Flexible Quantization and PEFT:
@@
library (seen in the Quant Config in Figure 1), which enables flexible fixed-point quantization. For most low-to-mid hardware setups, we recommend quantizing in 4-bit floating point via 4-bit NormalFloat (nf4). Additionally, Simul-LLM by default engages in PEFT Mangrulkar et al. (2022) via a passed configuration, although it does also support full model fine-tuning.
@@
object and a response template.
@@ 3.2 Evaluation Agent and Features
Support for Multiple Decoding Strategies: Variable latency constraints for possible inference environments demand flexibility in decoding strategies. As such, evaluation agents for Simul-LLM support several decoding strategies, including greedy and naive decoding, subword-based beam search for single-word decoding, and variations on Speculative Beam Search (SBS) (Zheng et al., 2019), including single word-based and chunk-wise SBS.
Scoring and Latency via SimulEval: SimulEval (Ma et al., 2020) is the premier SimulMT evaluation framework. Simul-LLM evaluation agents interface seamlessly with SimulEval, which handles the incremental source context and manages translation scoring and latency tracking regarding the translation hypothesis. In addition to the core functionality of SimulEval in terms of translation quality and latency scoring, Simul-LLM also features some extensions to SimulEval. This includes the capability to measure computationally aware text-based latency via the specification of an oracle, zero-latency transcription agent (effectively required for latency modeling, elaborated on in our Appendix) and support for COMET (Rei et al., 2020) translation quality evaluations (default model employed is COMET-DA (Rei et al., 2021)).
@@ 4 Adapting NMT LLMs to SimulMT
To intuit why this is an issue, suppose that an NMT LLM is accustomed to receiving the entire source sequence x before outputting a word or token . If engaging in a low-lagging wait-k where , then the output can now only be based on source context up to . Assuming the NMT LLM would typically rely on some source context where , then critical information is missing from its translation decision. We have conducted preliminary quantitative studies on this front by employing the proposed Simul-LLM framework, and these results are presented in Section 7.
@@ 5.1 Split Source-Target Prompt Structure
This lack of loss filtering can be especially problematic near the end of a given sequence’s translation. As a simple example, suppose that a given source sequence is of length and . In the first prompt structure where up to source words have been supplied and where , the LLM is effectively being fine-tuned for varying wait-k values ranging from wait-i to wait-k in a single example (as is predicted from to which is wait-k, is predicted from to which is wait-, and so on). If to the point where it is close to normal NMT levels of context, where , then the LLM is no longer being effectively fine-tuned with an appropriate amount of source context for a given translation (write) decision schedule.
@@ 5.2 Single Output Word Prompt Structure
The above problem can be entirely side-stepped by our proposed second approach, single output token prompt structure, that embeds only the current target translation hypothesis within the model output. As illustrated in the second prompt structure in Figure 3, instead of allowing target translation hypotheses from previous time-steps to be incorporated into the loss, the proposed prompt structure shifts those previous translation hypotheses into the prompt. Combined with the expanded examples that form a rigorous wait-k curriculum in terms of the fine-tuning dataset (rigorous meaning a complete curriculum as opposed to a random subset), inference behavior can be copied exactly for every fine-tuning example, completely closing the context mismatch between fine-tuning and inference.
@@ 6 Evaluation Methodology
To validate our proposed solutions to the aforementioned challenges and to test the capabilities of Simul-LLM as the first SimulMT LLM open-source framework, we engaged in several experiments allowing for comparisons among classical non-LLM-based NMT (Vaswani et al., 2017) / SimulMT architectures (Ma et al., 2019), NMT LLMs adapted for SimulMT, and SimulMT LLMs. All mentioned LLMs are fine-tuned Falcon-7B models, but Simul-LLM features an easy to extend framework and we support both Llama and Mistral-based models as well (results in Appendix).
NMT LLMs were fine-tuned for one epoch, as overfitting was observed beyond that point, which we intuit to be possibly due to the well-documented ability of LLMs to quickly memorize training sets (Biderman et al., 2023). In contrast, SimulMT LLMs were fine-tuned for either 2M random examples out of 5M examples (for wait-3 fine-tuning) or up to 5M examples (for wait-7 fine-tuning) on the expanded dataset due to computational constraints.
@@
SimulMT LLMs with Proposed Prompt | Wait-3 Fine-tuning LLM (2M samples) | 19.99 (3.41) | 23.68 (3.64) |
Wait-7 Fine-tuning LLM (2M samples) | 20.82 (3.44) | 25.18 (3.61) | |
Wait-7 Fine-tuning LLM (5M samples) | 21.86 (3.38) | 30.31 (3.34) | |
Wait-7 Fine-tuning LLM (2M samples, k=7) | 23.09 (6.71) | 28.92 (6.87) | |
Wait-7 Fine-tuning LLM (5M samples, k=7) | 24.93 (6.70) | 35.14 (6.73) |
@@ 6.1 Dataset Selection and Preprocessing
No standardized dataset exists for SimulMT with LLMs. Due to its popularity in speech-to-text simultaneous translation (SimulST), we employ MuST-C for our experiments333This allows for future work that explores multi-modal simultaneous LLMs, engaging in SimulST via a cascaded model structure with a transcription model for the source speech or a joint speech/text-to-text framework.. For the purposes of adapting MuST-C for text-to-text usage, we preprocess the dataset and filter out certain acoustic indicators (e.g., floating "-" characters representing pauses). In some cases, this resulted in significant changes to some samples of the test set, such as the removal of (Applause) acoustic indicators.
@@ 7.1 Exploration of Adapting NMT LLMs to SimulMT
In Table 1, we provide a breakdown of the performance of several different models, decoding strategies, and wait-k schedules. Regarding our exploration related to adapting NMT LLMs to SimulMT, we also include results related to our implementation of Speculative Beam Search (SBS) (Zheng et al., 2019). As demonstrated by these results, compared with classical models, LLMs fine-tuned for NMT are very capable of SimulMT upon being adapted during inference (even exceeding the score of the classical NMT transformer on en-de that performs non-simultaneous translation). It is worth reiterating that classical architectures often engage in subword-based wait-k whereas we employ word-based wait-k for LLMs, but the comparisons still serve as a useful reference.